NUCLEAR SPIN DEPENDENCE OF THE REACTION OF H_{3}^{+} WITH H_{2}

\underline{KYLE N. CRABTREE}, BRIAN A. TOM,^{a} CARRIE A. KAUFFMAN, BENJAMIN J. McCALL, Department of Chemistry, University of Illinois, Urbana, IL 61801.

The chemical reaction H_{3}^{+} + H_{2} → H_{2} + H_{3}^{+} is the simplest bimolecular reaction involving a polyatomic, and is possibly the most common such process occurring in the universe. Recent measurements of interstellar clouds have shown that the temperatures derived from the lowest rotational levels of H_{2} and H_{3}^{+} do not agree, and it is expected that this reaction plays a key role in this deviation. To investigate this process, we have measured the ortho/para ratio of H_{3}^{+} produced in this reaction by performing high resolution spectroscopy on its ν_{2} fundamental band in plasmas formed from various mixtures of ortho and para H_{2}. These measurments have been performed in a supersonic expansion discharge source and in a cooled hollow cathode cell to probe the reaction at a variety of temperatures at and below 300 K. Our results provide experimental evidence that the population distribution of the lowest levels of H_{3}^{+} is governed by the steady state of the H_{3}^{+} + H_{2} → H_{2} + H_{3}^{+} reaction, not by thermalization.

aCurrent Address: Department of Chemistry, United States Air Force Academy, CO 80840